### 27 Feb 2026 v1

# arXiv:2602.23780v1[eess.SP] 27 Feb 2026

Abstract
Convolution serves as a powerful operation for the regularization of functions. While polynomials inherently possess smoothness, it is particularly interesting to investigate their behavior under convolution. This interest stems from the fact that numerous engineering and physical phenomena can be modeled through such operations, including weighted averages, blurring effects and convolutional integral equations. In this work, we show that under certain mild conditions, the convolution with any even Schwartz function acts as an automorphism on the vector space of finite-order polynomials. We derive explicit equations for the inverse operation of this convolution, which are numerically simple to implement. In addition, we extend the deconvolution with (not necessarily even) Schwartz functions to a broader class of functions, including $L^{1}(\mathbb{R})$, $L^{2}(\mathbb{R})$, the Schwartz space and tempered distributions. Specifically, we establish a explicit rigorous formula for the deconvolution of a function or distribution that has been convolved with a Schwartz function, being a particular example the Weierstrass Transform. For the latter, we show that any Schwartz function and tempered distribution that has been transformed, can be recovered, in their respective topologies, by the limit of a sequence of linear combination of recursive convolutions. This provides a new formula for the inverse of the Weierstrass Transform that can be numerically implemented.
keywords:
Deconvolution, Weierstrass Transform, Polynomials, Schwartz Functions
pacs:
MSC Classification]44A35, 45E10, 42A85, 46F12, 44A15
1 Introduction
The Weierstrass transform with parameter $\varepsilon>0$ of a function $f:\mathbb{R}\to\mathbb{R}$, is given, when it exists, by
$F_{\varepsilon}(x):=(f*\varphi_{\varepsilon})(x)=\int_{\mathbb{R}}f(x-s)\varphi_{\varepsilon}(s)ds,$
where $x\mapsto\varphi_{\varepsilon}(x):=\frac{1}{\varepsilon\sqrt{4\pi}}e^{-\frac{x^{2}}{4\varepsilon^{2}}}$ is the so called Gaussian Kernel. The Gaussian Kernel is an even Schwartz function that satisfies the following properties. It has unit area, i.e., $\int_{\mathbb{R}}\varphi_{\varepsilon}=1$, it converges to the Dirac delta distribution as $\varepsilon\to 0^{+}$, and for any $f\in L^{p}_{loc}(\mathbb{R})$, $F_{\varepsilon}\in C^{\infty}(\mathbb{R})$ and $F_{\varepsilon}\to f$ as $\varepsilon\to 0^{+}$ in several types of convergence. In particular it $f\in L^{p}_{loc}(\mathbb{R})$ then $F_{\varepsilon}\to f$ as $\varepsilon\to 0^{+}$ pointwise and in $L^{p}$. It has been studied extensively in [widder1951necessary], where its inversion is presented as a differential operator, but also in [bilodeau1962weierstrass] where it is shown the relationship between the transform and the Hermite polynomials.
Observing the inversion formula in the aforementioned references, and how the Weierstrass transform acts on polynomials, it is natural to ask how the polynomials behave under a more general operation, and if there is an inverse for this operation which is simple to implement and compute That is, instead of focusing on the convolution with the Gaussian Kernel, we focus on the convolution with Schwartz convolution kernels. These questions, apart from its theoretical interest have interesting applications, since convolution with a function of the above characteristics can be thought as a weighted average operation, or equivalently, as a low pass filter, and so the process of deconvolution, i.e., the process of restoring an original signal from a filtered one, is of great application and theoretical interest. The theoretical importance of deconvolution is well-established. Classical treatments such as [berenstein1983some] and [hirschman2012convolution] provide systematic inversion formulas for several convolution transforms, [hohlfeld1993solution] provides explicit formulas for the solution of convolution integral equations by means of differentiation and [prost2003deconvolution] presents a formula for inverting the convolution when the convolution kernel has a Fourier transform with compact support. Moreover, in [baeumer2003inversion] is presented an inverse convolution approach by the limit of a finite sequence of linear combinations of exponential functions. More recently, in [kaiser2025deconvolutionarbitrarydistributionfunctions] deconvolution formulas for arbitrary distribution functions and densities using characteristic functions and geometric series methods are developed. Beyond theory, deconvolution has several practical applications and implementations. Numerical approaches are deeply treated in [Hansen2002-li], while [madden1996comparison] provides a comparative analysis of different deconvolution techniques. In [ulmer2010inverse] the deconvolution with Gaussian Kernels is treated and applied to radiotherapy problems, and in [fanton2021convolution] it is discussed several physical and engineering applications where (de)convolution is necessary. If the convolution kernel is unknown, blind deconvolution techniques become useful, as presented in [campisi2017blind] and [zhang2023fast]. Convolutions are also applied in more mathematical abstract situations, as in [cech2025optimalitymollifiers]. More recently, [Hohage_Marechal_Simar_Vanhems_2024] provides a novel framework for deconvolution techniques.
In this work we contribute both theoretically and practically to the deconvolution field with Theorem 2 and, particularly, with Theorem 3. In the former we present that the convolution with an even Schwartz function is an invertible operator on the space of finite order polynomials. In the latter we show that the inverse of the convolution with a Schwartz function can be explicitly computed as a limit of a sequence of linear combinations of recursive convolutions. More specifically, in Section 2 we show that, if the convolution kernel is an even Schwartz function, the convolution operator is an automorphism of polynomial vector spaces of finite degree. Moreover, we explicitly present the formula (Equations (7) and (8)) of the inverse of the convolution, and validate the results numerically. The inverse, curiously enough, consists on a linear combination of iterative convolutions of the convolved polynomial (see Theorem 2), from which it follows that any polynomial can be written as a linear combination of its iterative convolutions; that is, any polynomial is, in essence, a linear combination of its weighted averages, see Remark 4. We end the section discussing the main result (Theorem 3), which we consider the main contribution to the field of deconvolution, either theoretically and numerically speaking. It happens that, by passing to the limit in Theorem 2, we obtain Theorem 3, that is, we can recover as a limiting sequence of linear combinations of recursive convolutions the original function which has been convolved, obtaining, particularly, a new formula for the inverse of the Weierstrass Transform. For this special case of the Weierstrass transform, we also prove that any Schwartz function or tempered distribution can be recovered by passing to the limit in Equation (9), in their respective topologies. Clearly, several assumptions must be made clear for the original function and the convolution kernel, which do not contradict the traditional assumptions for the inverse of the convolution to exist in the Fourier transformed space. In Section 3 we present some applications of these results, validating our approach and providing some insight for new signal processing methods. We end this work with some conclusions in Section 4.
1.1 Notation
Before proceeding, it is important to introduce the notation used in this work. We denote by $\mathbb{N}=\{1,2,\dots\}$ the set of positive integers, and $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. All integrals must be thought with respect to the Lebesgue measure. Given a set $X$, we denote by $\operatorname{id}:X\to X$ the identity function defined as $x\in X\mapsto\operatorname{id}(x)=x$. The indicator function of a set $A$ is defined as
$\operatorname{ind}_{A}(x)=\begin{cases}1&x\in A,\\ 0&x\notin A\end{cases}.$
For a set $A$ we denote its closure as $\overline{A}$. The support of a real-valued function is defined as $\operatorname{supp}f:=\overline{\{x\in\operatorname{dom}f\mid f(x)\neq 0\}}$, where $\operatorname{dom}f$ is the domain of $f$. Given $x,y\in\mathbb{R}$ we define $xA+y=\{xa+y\mid x\in A\}$. If $n>2$ a natural number, by $\sum_{j\in 2\mathbb{N}_{0}}^{n}j$ we mean the sum of even nonnegative integers up to $n$ if $n$ is even, or up to n1n-1 if $n$ is odd. The sum $\sum_{j\in 2\mathbb{N}}^{n}j$ is then understood as before but for elements in $2\mathbb{N}$. Given two functions $f,g:\mathbb{R}^{d}\to\mathbb{R}$ with $d\in\mathbb{N}$ we denote, when it exists, its convolution as fgf*g. The space of Schwartz functions is denoted as $\mathscr{S}(\mathbb{R}^{d})$, while $\mathscr{S}_{M}(\mathbb{R}^{d})=\{f\in\mathscr{S}(\mathbb{R}^{d})\mid\int_{\mathbb{R}^{d}}f=1\}$, and $\mathscr{S}^{\prime}(\mathbb{R}^{d})$ denotes the space of tempered distributions. For $f\in\mathscr{S}^{\prime}(\mathbb{R}^{d})$ and $\psi\in\mathscr{S}(\mathbb{R}^{d})$ we denote the image of $\psi$ under $f$ either by $f(\psi)$ or $\langle f,\psi\rangle$. The space of $p\in[1,\infty]$ integrable functions as $L^{p}(\mathbb{R}^{d})$, and the space of locally integrable functions as $L^{p}_{loc}(\mathbb{R}^{d})$, i.e., $p$-integrable on each compact subset of $\mathbb{R}^{d}$. For a function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ we will denote by $\varphi_{\varepsilon}:=\frac{1}{\varepsilon^{d}}\varphi\circ\frac{\operatorname{id}}{\varepsilon}$ for any $\varepsilon>0$. We define the Fourier transform of a function $f:\mathbb{R}\to\mathbb{R}$, when it exists, by $\mathcal{F}(f)(\xi)=\hat{f}(\xi):=\int_{\mathbb{R}}f(x)e^{-2\pi i\xi x}dx$ for all $\xi\in\mathbb{R}$, and its inverse Fourier transform by $\mathcal{F}^{-1}(f)$. By LDCT we refer to the Lebesgue’s Dominated Convergence Theorem and by FFT we mean the Fast Fourier Transform.
2 Convolution and deconvolution of polynomials and arbitrary functions with Schwartz convolution kernels
Given $\varepsilon>0$ and even $\varphi\in\mathscr{S}(\mathbb{R})$, it is easy to check that if $p$ is an affine function, then $p*\varphi_{\varepsilon}=p$. This property does not hold in general for polynomials $p$ with a degree greater than one, that is, the fixed points for the (even) convolution operator are constants and affine functions. However, it is interesting to study if the convolution of a polynomial results in a polynomial, if this operation can be reversed, and if the inverse, if it exists, can be extended to more function spaces.
In this section we get the following results for a convolution kernel $\varphi\in\mathscr{S}(\mathbb{R})$ with $\int_{\mathbb{R}}\varphi\neq 0$ and any polynomial $p$ of degree $n\in\mathbb{N}_{0}$ defined in the real numbers.
  • If $\varphi$ is even, the convolution of $p$ with $\varphi$ results in a polynomial of the same degree (Theorem 1). In particular, if the original degree of the polynomial is $n\geq 2$, the convolution is the sum of the original polynomial with another polynomial of degree n2n-2, one or zero, depending on the parity of $n$ (Corollaries 1 and 2). In particular, the convolution with an even Schwartz function of a polynomial is another polynomial with the same asymptotic growth, and in addition the first and second leading coefficients are shared between $p$ and its convolution.
  • If $\varphi$ is even, for any $\varepsilon>0$, the convolution operator $p\mapsto p*\varphi_{\varepsilon}$ is an automorphism between the vector space of polynomials of the same degree, for any value of $\varepsilon>0$ (Theorem 2). This implies that the convolution with an even Schwartz kernel in the vector space of polynomials is invertible. The inverse can be computed as in Equations (7) and (8), and we provide an insight of its numerical simplicity in Corollary 3.
  • We prove in Theorem 3, that these results can be generalized by passing to the limit in Equation (7) of Theorem 2, for a certain class of functions, and without the need for the Schwartz convolution kernel to be even. This results are remarkable, since by an iterative process the original function can be restored via convolutions.
  • Some results can be extrapolated to polynomials in $\mathbb{R}^{d}$ with $d\geq 2$, see Section 2.3.
Refer to Section 3 to several applications of the previous results.
2.1 Polynomials in the real line
Let $p:\mathbb{R}\to\mathbb{R}$ be a polynomial of degree $n\in\mathbb{N}_{0}$, expressed in the form
$p(x)=a_{0}+a_{1}x+\dots+a_{n}x^{n}=\sum_{k=0}^{n}a_{k}x^{k},\quad(x\in\mathbb{R}),$ (1)
where $\{a_{k}\}_{k=0}^{n}\subset\mathbb{R},a_{n}\neq 0$. The vector space of all polynomials of degree $n\in\mathbb{N}_{0}$ will be denoted as $\mathscr{P}_{n}(\mathbb{R})$. We will say that $a_{n}$ is the leading coefficient, and $\{a_{n-k}\}_{k=0}^{r}$ are the $r$ leading coefficients for $0\leq r\leq n$. For a matter of convenience we will always exclude the case that $p\equiv 0$, i.e., the zero constant polynomial.
We will show that for any convolution kernel $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ it holds that $P_{\varepsilon}:=p*\varphi_{\varepsilon}$ is a polynomial of the same degree for $\varepsilon>0$. That is, no matter how $\varphi$ is constructed, that polynomials are, in some sense, invariant under the convolution with Schwartz kernels, i.e., the mapping $p\mapsto P_{\varepsilon}$ maps $\mathscr{P}_{n}(\mathbb{R})$ into $\mathscr{P}_{n}(\mathbb{R})$. But we will also prove that it is an automorphism if $\varphi$ is even. But first, we need a preliminary definition. We are going to restrict ourselves to the set $\mathscr{S}_{M}(\mathbb{R})$. But note that, being the convolution a linear mapping, and given $f\in\mathscr{S}(\mathbb{R})$, we can define $f/\int_{\mathbb{R}}f\in\mathscr{S}_{M}(\mathbb{R})$ as long as $\int_{\mathbb{R}}f\neq 0$. So we restrict ourselves from Schwartz functions with null integral.
Definition 1.
Let $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even. Define for each $m\in\mathbb{N}_{0}$ the constant
$c_{m}:=\int_{\mathbb{R}}x^{m}\varphi(x)dx.$ (2)
It is clear that $c_{0}=1$ and
$\int_{\mathbb{R}}x^{m}\varphi_{\varepsilon}(x)dx=\begin{cases}c_{m}\varepsilon^{m}&m\in 2\mathbb{N}_{0}\\ 0&m\in 2\mathbb{N}_{0}+1\end{cases}.$
Note that the integral is well defined since $\operatorname{id}^{m}\varphi\in\mathscr{S}(\mathbb{R})$ for any $m\in\mathbb{N}_{0}$.
Remark 1.
Note that if $p\in\mathscr{P}_{n}(\mathbb{R})$, then $\int_{\mathbb{R}}p\varphi_{\varepsilon}$ can be considered as a polynomial with even powers in $\varepsilon$ for any even $\varphi\in\mathscr{S}_{M}(\mathbb{R})$. Indeed if $n>2$ is even
$\varepsilon\mapsto h(\varepsilon):=\int_{\mathbb{R}}p(x)\varphi_{\varepsilon}(x)dx=\int_{\mathbb{R}}\sum_{k=0}^{n}a_{k}x^{k}\varphi_{\varepsilon}(x)dx=a_{0}+a_{2}c_{2}\varepsilon^{2}+\dots+a_{n}c_{n}\varepsilon^{n},$
while the result is clear if $n$ is odd. Moreover, it is clear that if $\varphi\in\mathscr{S}(\mathbb{R})$ is neither even nor odd, we can still define $c_{m}$ in the same way.
Theorem 1.
Let $p\in\mathscr{P}_{n}(\mathbb{R})$, $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and $\varepsilon>0$.
It holds that
$P_{\varepsilon}(x)=(p*\varphi_{\varepsilon})(x)=\sum_{j=0}^{n}\sum_{k\in 2\mathbb{N}_{0}}^{j}\binom{j}{k}a_{j}x^{j-k}c_{k}\varepsilon^{k},\quad(x\in\mathbb{R}).$ (3)
Therefore, the mapping $p\mapsto P_{\varepsilon}$ is a mapping of $\mathscr{P}_{n}(\mathbb{R})$ into itself.
Proof.
Let $m\in\mathbb{N}$. We know by the Binomial theorem that for all $x\in\mathbb{R}$,
$\int_{\mathbb{R}}(x-y)^{m}\varphi_{\varepsilon}(y)dy$ $=\int_{\mathbb{R}}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x^{m-k}y^{k}\varphi_{\varepsilon}(y)dy$
$=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x^{m-k}\int_{\mathbb{R}}y^{k}\varphi_{\varepsilon}(y)dy$
$=\sum_{k\in 2\mathbb{N}_{0}}^{m}\binom{m}{k}x^{m-k}c_{k}\varepsilon^{k},$
where the last equality comes from Definition 1 and the fact that $(-1)^{k}=1$ for even $k$. Now note that since the convolution is a bilinear operation
$P_{\varepsilon}(x)=\sum_{j=0}^{n}a_{j}\int_{\mathbb{R}}(x-y)^{j}\varphi_{\varepsilon}(y)dy=\sum_{j=0}^{n}a_{j}\sum_{k\in 2\mathbb{N}_{0}}^{j}\binom{j}{k}x^{j-k}c_{k}\varepsilon^{k},\quad(x\in\mathbb{R}),$
and since the sums are finite, we can use Fubini’s Theorem to end the proof of Equation (3). Note that the right hand side of Equation (3) is again a polynomial of at most degree $n$, proving the theorem. ∎
It is clear and immediate that if $\varphi$ is not even we still get a polynomial, the fact that we are restricting to even convolution kernels will become apparent in Theorem 2.
We now can improve upon the results of Theorem 1 using the following corollary.
Corollary 1.
Let $p\in\mathscr{P}_{n}(\mathbb{R})$, with $n\geq 2$ and $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and let $\varepsilon>0$. It holds that
$P_{\varepsilon}(x)=p(x)+\sum_{j=0}^{n-2}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right),\quad(x\in\mathbb{R}).$ (4)
That is, the convolution of a polynomial results in another polynomial with the same asymptotic growth, in fact the first two leading coefficients of $P_{\varepsilon}$ is the same as $p$.
Proof.
We proceed for the even case by induction over $n\geq 2$. Suppose first that $n=2$, then it is straightforward to see using Theorem 1 that
$P_{\varepsilon}(x)=a_{0}+a_{1}x+a_{2}x^{2}+a_{2}c_{2}\varepsilon^{2},\quad(x\in\mathbb{R}),$
and so Equation (4) is valid for $n=2$. Let’s suppose it is true for $n\in\mathbb{N}$ and proceed to prove that it is also true for $n+1$. Let $q\in\mathscr{P}_{n+1}(\mathbb{R})$, and denote by $r\in\mathscr{P}_{n}(\mathbb{R})$ the polynomial of degree $n$ obtained by making the leading coefficient of $q$ to be zero. Then by Theorem 1,
$Q_{\varepsilon}(x)=(q*\varphi_{\varepsilon})(x)=\sum_{j=0}^{n+1}\sum_{k\in 2\mathbb{N}_{0}}^{j}\binom{j}{k}a_{j}x^{j-k}c_{k}\varepsilon^{k}$
$=\sum_{j=0}^{n}\sum_{k\in 2\mathbb{N}_{0}}^{j}\binom{j}{k}a_{j}x^{j-k}c_{k}\varepsilon^{k}+\sum_{k\in 2\mathbb{N}_{0}}^{n+1}\binom{n+1}{k}a_{n+1}x^{n+1-k}c_{k}\varepsilon^{k}$
$=(r*\varphi_{\varepsilon})(x)+\sum_{k\in 2\mathbb{N}_{0}}^{n+1}\binom{n+1}{k}a_{n+1}x^{n+1-k}c_{k}\varepsilon^{k}$
$=\left[r(x)+\sum_{j=0}^{n-2}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right)\right]+\sum_{k\in 2\mathbb{N}_{0}}^{n+1}\binom{n+1}{k}a_{n+1}x^{n+1-k}c_{k}\varepsilon^{k}$
$=q(x)+\sum_{j=0}^{n-2}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right)+\sum_{k\in 2\mathbb{N}}^{n+1}\binom{n+1}{k}a_{n+1}x^{n+1-k}c_{k}\varepsilon^{k}.$
Suppose that $n$ is even, then
$\sum_{k\in 2\mathbb{N}}^{n+1}\binom{n+1}{k}a_{n+1}x^{n+1-k}c_{k}\varepsilon^{k}=x$ $^{n-1}\binom{n+1}{2}c_{2}\varepsilon^{2}a_{n+1}+x^{n-3}\binom{n+1}{4}c_{4}\varepsilon^{4}a_{n+1}$
$+\dots+x\binom{n+1}{n}c_{n}\varepsilon^{n}a_{n+1},$
which implies
$\sum_{j=0}^{n-2}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right)+\sum_{k\in 2\mathbb{N}}^{n+1}\binom{n+1}{k}a_{j}x^{n+1-k}c_{k}\varepsilon^{k}$
$=\sum_{j=0}^{n-1}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n+1-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right),$
completing the proof for even $n$. If $n$ is odd the proof is similar. ∎
Remark 2.
Given $p\in\mathscr{P}_{n}(\mathbb{R})$ and $n\geq 2$. We will denote by
$b_{j}:=\sum_{k\in 2\mathbb{N}}^{n-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j},\quad j\in\{0,1,\dots,n-2\},$
In doing so Equation (4) can be written as
$P_{\varepsilon}(x)=p(x)+\tilde{p}(x),\quad(x\in\mathbb{R}),$
where $\tilde{p}\in\mathscr{P}_{n-2}(\mathbb{R})$ with coefficients $\{b_{j}\}_{j=0}^{n-2}$, i.e., $\tilde{p}(x)=b_{0}+b_{1}x+\dots+b_{n-2}x^{n-2}$ for all $x\in\mathbb{R}$. Note that $b_{n-2}\neq 0$, since $b_{n-2}=c_{2}\varepsilon^{2}a_{n}\binom{n}{2}$ and $a_{n}\neq 0,c_{2}\neq 0$. This leads to the following corollary.
Corollary 2.
Let $p\in\mathscr{P}_{n}(\mathbb{R})$ with $n\geq 2$, $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and $\varepsilon>0$. Define by $T_{\varepsilon}:\mathscr{P}_{n}(\mathbb{R})\to\mathscr{P}_{n}(\mathbb{R})$ the map
$p\mapsto T_{\varepsilon}(p)=p*\varphi_{\varepsilon}.$
Given $k\in\mathbb{N}_{0}$ denote the composition of $T_{\varepsilon}$ with itself $k$ times as $T_{\varepsilon}^{k}$, i.e., the convolution applied recursively $k$ times, with $T_{\varepsilon}^{0}:=\operatorname{id}$. The following equation is true
$T_{\varepsilon}^{k}(p)=\sum_{j=0}^{k}\binom{k}{j}p_{j},$ (5)
where $p_{j}$ is defined recursively as follows. If $j=0$ then $p_{0}:=p$. If $j=1$ then $p_{j}:=\tilde{p}$ as in Remark 2, if $j=2$ then $p_{j}:=\tilde{\tilde{p}}$, which comes from $T(\tilde{p})=\tilde{p}+\tilde{\tilde{p}}$, as in Remark 2, and so on. Moreover the following statements are also true.
  1. 1.
    Suppose $n$ is even. If $k\in\mathbb{N}$ satisfies that $2k\leq n$ then $p_{j}\in\mathscr{P}_{n-2j}(\mathbb{R})$ for $j\in\{0,\dots,k\}$. If $2k>n$ then, $p_{j}\in\mathscr{P}_{n-2j}(\mathbb{R})$ for $j\in\{0,\dots,\frac{n}{2}-1\}$ and $p_{j}\in\mathscr{P}_{0}(\mathbb{R})$ for $j\in\{\frac{n}{2},\dots,k\}$, that is the last $k-\frac{n}{2}$ polynomials are constants.
  2. 2.
    Suppose $n$ is odd. If $k\in\mathbb{N}$ satisfies that $2k<n$ then $p_{j}\in\mathscr{P}_{n-2j}(\mathbb{R})$ for $j\in\{0,\dots,k\}\subset\{0,\dots,\frac{n-1}{2}\}$. If $2k>n$ then $p_{j}\in\mathscr{P}_{n-2j}(\mathbb{R})$ for $j\in\{0,\dots,\frac{n-1}{2}-1\}$ and $p_{j}\in\mathscr{P}_{1}(\mathbb{R})$ for $j\in\{\frac{n-1}{2},\dots,k\}$, that is, the last $k-\frac{n-1}{2}$ polynomials are affine functions.
Proof.
We first proceed with the proof of Equation (5). Fix $n\in\mathbb{N}$ with $n\geq 2$. We proceed by induction over $k$. For $k=1$ it holds that
$T_{\varepsilon}(p)=P_{\varepsilon}=p+\tilde{p}=p+p_{1},$
where $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ by Corollary 1. Note that $T$ is linear, which implies
$T_{\varepsilon}(T_{\varepsilon}(p))=T_{\varepsilon}(p+p_{1})=T_{\varepsilon}(p)+T_{\varepsilon}(p_{1})=p+p_{1}+p_{1}+p_{2}=p+2p_{1}+p_{2}.$
So the statement is proven for $k=2$, because $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ by Corollary 1 and Remark 2, which by the same conclusions imply that $p_{2}\in\mathscr{P}_{\max\{n-4,0\}}(\mathbb{R})$.
Suppose the statement is true for $k\in\mathbb{N}$, and $2k\leq n$, that is $T_{\varepsilon}^{k}(p)=\sum_{j=0}^{k}\binom{k}{j}p_{j}$, with $p_{j}\in\mathscr{P}_{n-2j}(\mathbb{R})$ with $j\in\{0,\dots,k\}$. Note $T_{\varepsilon}$ is linear by the properties of the convolution, thus
$T_{\varepsilon}^{k+1}(p)=T_{\varepsilon}(T_{\varepsilon}^{k}(p))=T_{\varepsilon}\left(\sum_{j=0}^{k}\binom{k}{j}p_{j}\right)=\sum_{j=0}^{k}\binom{k}{j}T_{\varepsilon}(p_{j})=\sum_{j=0}^{k}\binom{k}{j}(p_{j}+p_{j+1}),$
and then
$T_{\varepsilon}^{k+1}(p)$ $=\sum_{j=1}^{k}\left(\binom{k}{j}+\binom{k}{j-1}\right)p_{j}+p_{0}+p_{k+1}$
$=\sum_{j=1}^{k}\binom{k+1}{j}p_{j}+p_{0}+p_{k+1}=\sum_{j=0}^{k+1}\binom{k+1}{j}p_{j},$
where we used the fact that $\binom{k}{j}+\binom{k}{j-1}=\binom{k+1}{j}$. Moreover, $p_{k}\in\mathscr{P}_{n-2k}(\mathbb{R})$, which implies by Corollary 1 and Remark 2 that $p_{k+1}\in\mathscr{P}_{n-2(k+1)}(\mathbb{R})$. This proves Equation (5). Note that the formula is always valid, independently if $n$ is even or odd, so what we need now is to prove it for when $2k>n$.
  1. 1.
    Suppose $n$ is even and $2k\leq n$, then the previous discussion is valid, and so $p_{j}\in\mathscr{P}_{n-2j}(\mathbb{R})$ for $j\in\{0,\dots,k\}$. Suppose that $2k>n$, then $k>\frac{n}{2}$ and for $j=n/2$ it holds by the previous proof that $p_{j}\in\mathscr{P}_{0}(\mathbb{R})$, i.e., it is a constant. Since constant functions are invariant under convolution with $\varphi\in\mathscr{S}_{M}(\mathbb{R})$, it holds that $T_{\varepsilon}(p_{j})=p_{j}$ for $j\in\{\frac{n}{2},\dots,k\}$, and this proves the first assertion.
  2. 2.
    Suppose $n$ is odd and $2k<n$, then $2k\leq n-1$, so the previous discussion is valid, that is $p_{j}\in\mathscr{P}_{n-2j}(\mathbb{R})$ for $j\in\{0,\dots,k\}$. Now, if $2k>n$ then $2k\geq n+1$, and for $j=\frac{n-1}{2}$ it holds that $p_{j}\in\mathscr{P}_{n-2j}(\mathbb{R})=\mathscr{P}_{1}(\mathbb{R})$, but then $p_{j}$ is affine, which implies that $T_{\varepsilon}(p_{j})=p_{j}$ for $j=\frac{n-1}{2}$ and so for $j\in\{\frac{n-1}{2},\dots,k\}$, proving the second assertion.
It will be useful to compute the polynomials $p_{j}$ in Corollary 2 as linear combinations of repeated convolutions of $p_{0}=p$, i.e., of the original polynomial. This will allow to obtain a formula for the inversion of the convolution on the space $\mathscr{P}_{n}(\mathbb{R})$. We prove that this is possible in the next lemma.
Lemma 1.
Let $p\in\mathscr{P}_{n}(\mathbb{R})$ with $n\geq 2$, $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and $\varepsilon>0$. Defining $T_{\varepsilon}:\mathscr{P}_{n}(\mathbb{R})\to\mathscr{P}_{n}(\mathbb{R})$ as in Corollary 2 as well as $p_{j}$ for $j\in\mathbb{N}_{0}$, then
$p_{j}=\sum_{k=0}^{j}\binom{j}{k}(-1)^{k}T_{\varepsilon}^{j-k}(p),$ (6)
where $T_{\varepsilon}^{0}:=\operatorname{id}$ for all $\varepsilon>0$ and $p_{0}:=p$.
Proof.
We proceed by induction over $j$. By definition it is true for $j=0$, and for $j=1$ we know by Theorem 1 and Corollary 2 that $T_{\varepsilon}(p_{0})=p_{0}+p_{1}$ and therefore $p_{1}=T_{\varepsilon}(p_{0})-p_{0}$ for which Equation (6) clearly holds. Suppose now that it is true for $j\in\mathbb{N}$. Then $T_{\varepsilon}(p_{j})=p_{j}+p_{j+1}$ and so
$p_{j+1}=T_{\varepsilon}(p_{j})-p_{j}$ $=\sum_{k=0}^{j}\binom{j}{k}(-1)^{k}T_{\varepsilon}^{j+1-k}(p_{0})+\sum_{k=0}^{j}\binom{j}{k}(-1)^{k+1}T_{\varepsilon}^{j-k}(p_{0})$
$=\sum_{k=0}^{j}\binom{j}{k}(-1)^{k}T_{\varepsilon}^{j+1-k}(p_{0})+\sum_{k=1}^{j+1}\binom{j}{k-1}(-1)^{k}T_{\varepsilon}^{j+1-k}(p_{0})$
$=T_{\varepsilon}^{j+1}(p_{0})+(-1)^{j+1}T_{\varepsilon}(p_{0})+\sum_{k=1}^{j}\binom{j+1}{k}(-1)^{k}T_{\varepsilon}^{j+1-k}(p_{0})$
$=\sum_{k=0}^{j+1}\binom{j+1}{k}(-1)^{k}T_{\varepsilon}^{j+1-k}(p_{0}),$
where the fact $\binom{j+1}{k}=\binom{j}{k-1}+\binom{j}{k}$ has been used in the second to last inequality. This proves Equation (6), thus the lemma. ∎
We now prove that the mapping $T_{\varepsilon}$ defined in Corollary 2, is an isomorphism of the vector space $\mathscr{P}_{n}(\mathbb{R})$ onto itself. For this purpose we will use the results from Corollary 1 and Remark 2 in a recursive manner. This result has some remarkable properties: in $\mathscr{P}_{n}(\mathbb{R})$ the convolution with an even convolution Schwartz kernel is always a reversible operation. Therefore, we can consider this to be an extension of the Weierstrass Transform of polynomials, in which $\varphi$ is the Gaussian Kernel, and so $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ is even. See [bilodeau1962weierstrass] for a discussion of Hermite polynomials and the Weierstrass function. Finally, note in the proof of the following Theorem that the even part plays a crucial role. Indeed, the even part accounts for two reasons. First, if $p\in\mathscr{P}_{n}(\mathbb{R})$ then $T_{\varepsilon}(p)=p+p_{1}$ with $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$, that is, it reduces $p_{1}$ two degrees below $p$. Also if $p$ is an affine or constant function then $T_{\varepsilon}(p)=p$. This is what will allow us to prove the automorphism and the computation of the inverse.
Theorem 2.
Let $n\in\mathbb{N}$, $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and $\varepsilon>0$. Then $T_{\varepsilon}:\mathscr{P}_{n}(\mathbb{R})\to\mathscr{P}_{n}(\mathbb{R})$, defined as in Corollary 2 is an automorphism. Moreover, given $p\in\mathscr{P}_{n}(\mathbb{R})$ with $n$ even
$T_{\varepsilon}^{-1}(p)=\sum_{j=0}^{n/2}(-1)^{j}p_{j}=\sum_{j=0}^{n/2}\sum_{k=0}^{j}(-1)^{j+k}\binom{j}{k}T_{\varepsilon}^{j-k}(p)=\sum_{j=0}^{n/2}(-1)^{n/2-j}\binom{n/2+1}{j}T_{\varepsilon}^{n/2-j}(p),$ (7)
where $p_{0}:=p$, and $T_{\varepsilon}^{0}:=\operatorname{id}$, while if $n$ is odd
$T_{\varepsilon}^{-1}(p)=\sum_{j=0}^{(n-1)/2}(-1)^{j}p_{j}=\sum_{j=0}^{(n-1)/2}(-1)^{\frac{n-1}{2}-j}\binom{\frac{n+1}{2}}{j}T_{\varepsilon}^{\frac{n-1}{2}-j}(p).$ (8)
Proof.
The fact that $T_{\varepsilon}$ is linear follows directly from the definition of the convolution. We proceed to prove that it is bijective. Suppose first that $n=2$. We know by Corollary 2 that $T_{\varepsilon}(p)=p+p_{1}$ where $p_{1}\in\mathscr{P}_{0}(\mathbb{R})$, and therefore $T_{\varepsilon}(p_{1})=p_{1}$. If $p,q\in\mathscr{P}_{2}(\mathbb{R})$ such that $p(x)=a_{0}+a_{1}x+a_{2}x^{2}$ and $q(x)=b_{0}+b_{1}x+b_{2}x^{2}$ for all $x\in\mathbb{R}$, then $p_{1}(x)=a_{2}\varepsilon^{2}c_{2}$ and $q_{1}(x)=b_{2}\varepsilon^{2}c_{2}$. Thus
$T_{\varepsilon}(p)=T_{\varepsilon}(q)\iff p+p_{1}=q+q_{1},$
but since $p+p_{1}\in\mathscr{P}_{2}(\mathbb{R})$ and $q+q_{1}\in\mathscr{P}_{2}(\mathbb{R})$ with the same leading coefficients as $p$ and $q$ by Corollary 1, and two polynomials of the same degree are equal if and only if all of its coefficients are equal, it follows that
$b_{2}=a_{2},b_{1}=a_{1},b_{0}+b_{2}\varepsilon^{2}c_{2}=a_{0}+a_{2}\varepsilon^{2}c_{2},$
from which it follows that $b_{0}=a_{0}$, that is $p=q$, proving that $T_{\varepsilon}$ is injective. Clearly given $p\in\mathscr{P}_{2}(\mathbb{R})$ the polynomial $p-p_{1}$ satisfies that $T_{\varepsilon}(p-p_{1})=T_{\varepsilon}(p)-T_{\varepsilon}(p_{1})=p$ proving that $T_{\varepsilon}$ is an isomorphism of $\mathscr{P}_{2}(\mathbb{R})$ onto itself. Therefore, we can know generalize this concepts by induction.
  • $T_{\varepsilon}$ is injective. Let $n\in\mathbb{N}$. Suppose $T(p)=T(q)$, then $p+p_{1}=q+q_{1}$. Since $p_{1},q_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ this implies that $a_{n}=b_{n}$ and $a_{n-1}=b_{n-1}$. Now, note that by Remark 2 that
    $a_{n-2}+\binom{n}{2}c_{2}\varepsilon^{2}a_{n}=b_{n-2}+\binom{n}{2}c_{2}\varepsilon^{2}b_{n},$
    and therefore $a_{n-2}=b_{n-2}$. We can continue in this manner to prove that $a_{j}=b_{j}$ for all $j\in\{0,\dots,n\}$, that is, that $p=q$, proving the injection.
  • $T_{\varepsilon}$ is surjective. Suppose $n$ is even. If $n=2$, by the previous proof, for all $p\in\mathscr{P}_{2}(\mathbb{R})$ there exists $q_{p}\in\mathscr{P}_{2}(\mathbb{R})$ such that $T_{\varepsilon}(q_{p})=p$. Now let $p\in\mathscr{P}_{4}(\mathbb{R})$, we know that $T_{\varepsilon}(p)=p+p_{1}$ with $p_{1}\in\mathscr{P}_{2}(\mathbb{R})$, and so there exists $q_{p_{1}}\in\mathscr{P}_{2}(\mathbb{R})$ such that $T_{\varepsilon}(q_{p_{1}})=p_{1}$. Therefore $T_{\varepsilon}(p-q_{p_{1}})=T_{\varepsilon}(p)-T_{\varepsilon}(q_{p_{1}})=p+p_{1}-p_{1}=p$, and so $T_{\varepsilon}$ is a surjection of $\mathscr{P}_{4}(\mathbb{R})$ onto itself. We can now easily proceed by induction to prove it is a surjection for $n\in 2\mathbb{N}$.
    Suppose now that $n=3$. Then $T(p)=p+p_{1}$ with $p\in\mathscr{P}_{1}(\mathbb{R})$, which implies that $T(p_{1})=p_{1}$. Therefore, $q=p-p_{1}$ satisfies $T_{\varepsilon}(q)=T_{\varepsilon}(p)-T_{\varepsilon}(p_{1})=p$ proving the surjection for $n=3$. Suppose it is true for $3\leq n\in 2\mathbb{N}+1$. Given $p\in\mathscr{P}_{n+2}(\mathbb{R})$ we know that $T_{\varepsilon}(p)=p+p_{1}$ with $p_{1}\in\mathscr{P}_{n}(\mathbb{R})$, and so by the induction hypothesis there exists $q_{1}\in\mathscr{P}_{n}(\mathbb{R})$ such that $T_{\varepsilon}(q_{1})=p_{1}$. Therefore, choosing $q=p-q_{1}$ the surjection is proven.
We proceed now to prove Equation (7). For $n=2$ the formula reads $T_{\varepsilon}^{-1}(p)=p-p_{1}$ which is true, since $T_{\varepsilon}(T_{\varepsilon}^{-1}(p))=T_{\varepsilon}(p)-T_{\varepsilon}(p_{1})=p+p_{1}-p_{1}=p$. Suppose then it is true for $n\geq 2$. Given $p\in\mathscr{P}_{n+2}(\mathbb{R})$ we know that $T_{\varepsilon}(p)=p+p_{1}$ where $p_{1}\in\mathscr{P}_{n}(\mathbb{R})$. If $q=p-T_{\varepsilon}^{-1}(p_{1})$, then $T_{\varepsilon}(q)=T_{\varepsilon}(p)-p_{1}=p+p_{1}-p_{1}=p$. That is,
$q=T_{\varepsilon}^{-1}(p)=p-T_{\varepsilon}^{-1}(p_{1})=p-\sum_{j=0}^{n/2}(-1)^{j}p_{j+1}=p_{0}+\sum_{j=1}^{n/2+1}(-1)^{j}p_{j}=\sum_{j=0}^{(n+2)/2}(-1)^{j}p_{j},$
proving the first equality of Equation (7). The second equality comes from Lemma 1. Finally, the third equality can we proven by induction. For $n=2$ we have that $T_{\varepsilon}^{-1}(p)=2p-T_{\varepsilon}(p)$ for both formulas. Suppose then it holds for $n\geq 2$ with $n$ even. Then for $n+2$
$T_{\varepsilon}^{-1}(p)$ $=\sum_{j=0}^{(n+2)/2}\sum_{k=0}^{j}(-1)^{j+k}\binom{j}{k}T_{\varepsilon}^{j-k}(p)$
$=\sum_{j=0}^{n/2}\sum_{k=0}^{j}(-1)^{j+k}\binom{j}{k}T_{\varepsilon}^{j-k}(p)+\sum_{k=0}^{n/2+1}(-1)^{n/2+1+k}\binom{n/2+1}{k}T_{\varepsilon}^{n/2+1-k}(p)$
$=\sum_{j=0}^{n/2}(-1)^{\frac{n}{2}-j}\binom{\frac{n}{2}+1}{j}T_{\varepsilon}^{\frac{n}{2}-j}(p)+\sum_{k=1}^{\frac{n}{2}+1}(-1)^{\frac{n}{2}+1-k}\binom{\frac{n}{2}+1}{k}T_{\varepsilon}^{\frac{n}{2}+1-k}(p)$
$+(-1)^{\frac{n}{2}+1}T_{\varepsilon}^{\frac{n}{2}+1}(p)$
$=\sum_{j=0}^{n/2}(-1)^{n/2-j}\left[\binom{\frac{n}{2}+1}{j}+\binom{\frac{n}{2}+1}{j+1}\right]T_{\varepsilon}^{\frac{n}{2}-j}(p)+(-1)^{\frac{n}{2}+1}\binom{\frac{n+2}{2}+1}{0}T_{\varepsilon}^{\frac{n}{2}+1}(p)$
$=\sum_{j=0}^{(n+2)/2}(-1)^{\frac{n+2}{2}-j}\binom{\frac{n+2}{2}+1}{j}T_{\varepsilon}^{\frac{n+2}{2}-j}(p).$
proving Equation (7). The proof for the case where $n$ odd, i.e., Equation (8), is identical to the even case. ∎
Corollary 3.
Equation (7) can also be computed as
$T_{\varepsilon}^{-1}=\sum_{j=0}^{n/2}(-1)^{j}\binom{n/2+1}{j+1}T_{\varepsilon}^{j}$
which in turn implies that in $\mathscr{P}_{n}(\mathbb{R})$, with even $n$,
$T_{\varepsilon}^{-1}\circ T_{\varepsilon}=\operatorname{id}=-\sum_{j=1}^{n/2+1}\binom{n/2+1}{j}(-T_{\varepsilon})^{j}=\operatorname{id}-(\operatorname{id}-T_{\varepsilon})^{n/2+1},$
that is $(\operatorname{id}-T_{\varepsilon})^{n/2+1}\equiv 0$ on $\mathscr{P}_{n}(\mathbb{R})$.
Proof.
The first result follows from a change of indexes in the sum. The second follows noting that, if $f:V\to V$ is a linear mapping, where $V$ is any vector space, then, by the binomial theorem for linear functions, $\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}f^{k}=(\operatorname{id}-f)^{n}$ where $\operatorname{id}:V\to V$ is the identity on $V$. ∎
Note from Corollary 3 it is only needed to compute $n/2$ convolutions numerically speaking, because each time the index of the sum advances, it is only needed to convolve the previously convolved function.
Remark 3.
Theorem 2 states that, for polynomials, convolution is invertible as long as the kernel is an even (Schwartz) function. Note that convolution is, practically speaking, the filtering of the function with a smooth filter in $\mathscr{S}_{M}(\mathbb{R})$. Note that since $\mathscr{S}_{M}(\mathbb{R})\subset\mathscr{S}(\mathbb{R})\subset L^{2}(\mathbb{R})$ the filter has finite energy. What the Theorem states is that, even after having (for example) low-filtered a polynomial via convolution with an even convolution kernel (e.g., Gaussian Kernel, or the compactly supported convolution kernel $x\mapsto\varphi(x)=\operatorname{ind}_{(-1,1)}\exp(-1/(1-x^{2}))$), we can recover the original function uniquely. This powerful fact can also be numerically implemented easily by Equations (7) and (8). See Section 3 for an application related to this remark.
Remark 4.
Suppose $n$ is even. Note that $T_{\varepsilon}(T^{-1}_{\varepsilon}(p))=p$ and $T_{\varepsilon}$ is linear, which implies
$p=\sum_{j=0}^{n/2}(-1)^{n/2-j}\binom{n/2+1}{j}T_{\varepsilon}^{n/2+1-j}(p),$
that is, any polynomial can be expressed as a linear combination of its convolutions with even Schwartz convolution kernels, no matter the $\varepsilon>0$ nor the (even) convolution kernel. This is an interesting result: the sum of the convolved versions of the polynomial generate the same polynomial. For example, for $n=4$,
$p=T_{\varepsilon}^{3}(p)-3T_{\varepsilon}^{2}(p)+3T_{\varepsilon}(p)=p*(3\varphi_{\varepsilon}-3\varphi_{\varepsilon}*\varphi_{\varepsilon}+\varphi_{\varepsilon}*\varphi_{\varepsilon}*\varphi_{\varepsilon}).$
An in general, for even $n$ we have that $\sum_{j=0}^{n/2}(-1)^{n/2-j}\binom{n/2+1}{j}T_{\varepsilon}^{n/2-j}(\varphi_{\varepsilon})$ is an identity operator for the convolution of polynomials in $\mathscr{P}_{n}(\mathbb{R})$ with even Schwartz kernels. Clearly, the identity operator is not a polynomial, and thus they do not form a group. Moreover, note that being $T_{\varepsilon}$ an automorphism in $\mathscr{P}_{n}(\mathbb{R})$ then it cannot hold that $T_{\varepsilon}(p)$ is identically zero for any $p\in\mathscr{P}_{n}(\mathbb{R})$ except for $p\equiv 0$. This is congruent with information theory, we can always recover the original polynomial after convolution because there is always information about it after the convolution, i.e., $T_{\varepsilon}(p)\not\equiv 0$ for any $p\not\equiv 0$.
We end this section by noting the following. There is nothing special about $\varphi$ being a Schwartz function, since as it can be seen in the proofs of the previous results, the rapidly decreasing property of $\varphi$ is only needed so that the convolution is well posed. Therefore, any other even function for which $p*\varphi$, as well as its recursive convolutions, make sense, is equally valid. Nevertheless, in the following section we will show that the requirement for $\varphi$ to be a Schwartz function is essential to carry the generalization to more function spaces.
2.2 The extension to a bigger space of functions: The deconvolution of arbitrary functions
The natural question now is to ask if, for a given function $f$, it holds (under certain assumptions about $f$) that, by passing to the limit in Equation (7), we can recover $f$ from $T_{\varepsilon}(f)$. That is, if $T_{\varepsilon,n}^{-1}$ is the inverse of the automorphism $T_{\varepsilon}$ in $\mathscr{P}_{n}(\mathbb{R})$ for some even $\varphi\in\mathscr{S}(\mathbb{R})$ (see Theorem 2) does it hold that
$\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))=f,$
in some sense of convergence? The answer is affirmative under some conditions on the Fourier transform of $\varphi\in\mathscr{S}$ and different assumptions for $f$. We prove these results in the next theorem, in which the assumption that $\varphi$ is even is not always required. As mention in the introduction, we will denote $\mathcal{F}\left(f\right)$ or $\widehat{f}$ the Fourier transform of a function $f$ whenever it exists, and as $\mathcal{F}$ the Fourier transform as an operator. For its properties refer to [grubb2008distributions] and [grafakos2008classical].
Theorem 3.
Let $\varphi\in\mathscr{S}(\mathbb{R})$ and let $\varepsilon>0$. Suppose that for all $\xi\in\mathbb{R}$ it holds that $0<\widehat{\varphi_{\varepsilon}}(\xi)<2$. Let the operator $f\mapsto T_{\varepsilon}(f)=f*\varphi_{\varepsilon}$, when it makes sense for $f$, and for $n\in\mathbb{N}$ let $T_{\varepsilon,n}^{-1}$ be the operator
$f\mapsto T_{\varepsilon,n}^{-1}(f)=\sum_{k=0}^{n}(-1)^{k}\binom{n+1}{k+1}T_{\varepsilon}^{k}(f),$ (9)
when it makes sense for $f$, where $T_{\varepsilon}^{k}$ is the $k$-th folding composition of $T_{\varepsilon}$. That is, the convolution applied repetitively $k$ times. If $f\in\mathscr{S}^{\prime}(\mathbb{R})$ denote by $f*\varphi_{\varepsilon}$ the tempered distribution $\langle f*\varphi_{\varepsilon},\psi\rangle=\langle f,\check{\varphi_{\varepsilon}}*\psi\rangle$ for all $\psi\in\mathscr{S}(\mathbb{R})$, where $\check{\varphi_{\varepsilon}}(x)=\varphi_{\varepsilon}(-x)$ for all $x\in\mathbb{R}$. Extend the definition of the operators $T_{\varepsilon}$ and $T_{\varepsilon,n}^{-1}$ to $\mathscr{S}^{\prime}$. Then the following statements are true
  1. 1.
    If $f\in L^{2}(\mathbb{R})$ then $\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))=f$ in $L^{2}(\mathbb{R})$.
  2. 2.
    If $f\in L^{1}(\mathbb{R})$ with $\widehat{f}\in L^{1}(\mathbb{R})$ then $\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))=f$ pointwise.
  3. 3.
    Let $\mathds{1}\in\mathscr{S}^{\prime}$ be the one tempered distribution (i.e., $\langle\mathds{1},\psi\rangle=\int_{\mathbb{R}}\psi$ for all $\psi\in\mathscr{S}(\mathbb{R}))$, let $\delta\in\mathscr{S}^{\prime}$ be the Dirac delta distribution, let $\mathrm{pv}\frac{1}{x}\in\mathscr{S}^{\prime}$ be the Cauchy Principal Value distribution, and for $y\in\mathbb{R}$ let $e^{2\pi iy\operatorname{id}}$ the distribution that arises from the (slowly increasing) function $x\mapsto e^{2\pi ixy}$. Then, for any of these distributions $u\in\mathscr{S}^{\prime}$ it holds that $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(u))\to u$ as $n\to\infty$ in the distributional topology. The same holds for any linear combination or/and translations of these distributions.
  4. 4.
    Suppose now that $\varphi$ is the Gaussian Kernel, i.e., $\varphi(x)=\frac{1}{\sqrt{4\pi}}e^{-(x/2)^{2}}$ for all $x\in\mathbb{R}$. Then, given $f\in\mathscr{S}(\mathbb{R})$ it holds that $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\to f$ as $n\to\infty$ in the Schwartz topology, and for any $f\in\mathscr{S}^{\prime}(\mathbb{R})$, it holds $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\to f$ as $n\to\infty$ in the tempered distribution sense.
Proof.
We prove each statement separately.
  1. 1.
    First note that since $f\in L^{2}(\mathbb{R})$ and $\varphi_{\varepsilon}\in\mathscr{S}(\mathbb{R})$ then $T_{\varepsilon}(f)=f*\varphi_{\varepsilon}\in L^{2}(\mathbb{R})$ by Young’s inequality for convolution [folland1999real], which implies that $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\in L^{2}(\mathbb{R})$ and so the operator makes sense. By the Fourier transform properties, it holds that $\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)\in L^{2}(\mathbb{R})$ for all $n\in\mathbb{N}$ and all $f\in L^{2}(\mathbb{R})$. Moreover $\widehat{\varphi_{\varepsilon}}\in\mathscr{S}(\mathbb{R})$, and since $\mathcal{F}\left(T_{\varepsilon}(f)\right)=\widehat{f}\widehat{\varphi_{\varepsilon}}$, by induction it is easily seen that $\mathcal{F}\left(T^{k}_{\varepsilon}(f)\right)=\widehat{f}(\widehat{\varphi_{\varepsilon}}\dots\widehat{\varphi_{\varepsilon}})=\widehat{f}\widehat{\varphi_{\varepsilon}}^{k}$ for all $k\in\mathbb{N}$. Thus, by the linearity of the Fourier transform
    $\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)=\sum_{k=0}^{n}(-1)^{k}\binom{n+1}{k+1}\widehat{f}\widehat{\varphi_{\varepsilon}}^{k+1}=\hat{f}\sum_{k=0}^{n}(-1)^{k}\binom{n+1}{k+1}\widehat{\varphi_{\varepsilon}}^{k+1}.$
    Moreover,
    $\sum_{k=0}^{n}(-1)^{k}\binom{n+1}{k+1}\widehat{\varphi_{\varepsilon}}^{k+1}=\sum_{k=1}^{n+1}(-1)^{k-1}\binom{n+1}{k}\widehat{\varphi_{\varepsilon}}^{k}=-\sum_{k=1}^{n+1}\binom{n+1}{k}\left(-\widehat{\varphi_{\varepsilon}}\right)^{k},$
    but by the binomial theorem $(1-x)^{n+1}=\sum_{k=0}^{n+1}\binom{n+1}{k}(-x)^{k}$, and therefore
    $-\sum_{k=1}^{n+1}\binom{n+1}{k}\left(-\widehat{\varphi_{\varepsilon}}\right)^{k}=1-(1-\widehat{\varphi_{\varepsilon}})^{n+1}.$
    Since $\widehat{\varphi_{\varepsilon}}\in\mathscr{S}$ then so it does $\widehat{\varphi_{\varepsilon}}^{k}$ for all $k\in\mathbb{N}$, and so $1-(1-\widehat{\varphi})^{n+1}\in\mathscr{S}$ for all $n\in\mathbb{N}$, being $\mathscr{S}$ a vector space. Now we claim that $\widehat{f}(1-(1-\widehat{\varphi_{\varepsilon}})^{n})\to\widehat{f}$ in $L^{2}(\mathbb{R})$. Indeed, note that $|\widehat{f}|^{2}\in L^{1}(\mathbb{R})$, and also
    $|\widehat{f}(1-(1-\widehat{\varphi_{\varepsilon}})^{n})-\widehat{f}|^{2}=|\widehat{f}|^{2}|1-\widehat{\varphi_{\varepsilon}}|^{2n}<|\widehat{f}|^{2},\quad\forall n\in\mathbb{N},$
    due to the fact that $0<\widehat{\varphi_{\varepsilon}}<2$. It also holds that [^1] If there exists some $x\in\mathbb{R}$ in which $\widehat{\varphi_{\varepsilon}}(x)=0$, then we cannot conclude pointwise convergence. This subtle assumption has important consequences. Indeed, if the frequency content has been removed, it cannot be recovered uniquely, since there could be an infinite amount of functions for which that frequency is also zero. However, we could also consider pointwise almost everywhere convergence, for which the proof is identical. $(1-\widehat{\varphi_{\varepsilon}})^{n}\to 0$ pointwise as $n\to\infty$. Thus, applying the Lebesgue’s Dominated Convergence Theorem (LDCT)
    $\lim_{n\to\infty}\int_{\mathbb{R}}|\widehat{f}(x)(1-(1-\widehat{\varphi_{\varepsilon}}(x))^{n}-\widehat{f}(x)|^{2}dx=\lim_{n\to\infty}\int_{\mathbb{R}}|\widehat{f}(x)|^{2}|1-\widehat{\varphi_{\varepsilon}}(x)|^{2n}dx=0.$
    Thus, $\lim_{n\to\infty}\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)=\widehat{f}$ in $L^{2}(\mathbb{R})$. Since the Fourier transform is an isometric isomorphism in $L^{2}(\mathbb{R})$, it holds that
[…]
    $=\lim_{n\to\infty}\mathcal{F}^{-1}\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)=\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)),$
    in $L^{2}(\mathbb{R})$. This proves the statement.
  2. 2.
    We know by the previous statement that
    $\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)=\left(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1}\right)\widehat{f},\quad\forall n\in\mathbb{N}.$
[…]
    $\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))(x)$ $=\lim_{n\to\infty}\mathcal{F}^{-1}\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)(x)$
    $=\lim_{n\to\infty}\int_{\mathbb{R}}\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)(\xi)e^{2\pi i\xi x}d\xi$
    $=\lim_{n\to\infty}\int_{\mathbb{R}}\left(1-(1-\widehat{\varphi_{\varepsilon}}(\xi))^{n+1}\right)\widehat{f}(\xi)e^{2\pi i\xi x}d\xi$
    $=\int_{\mathbb{R}}\widehat{f}(\xi)e^{2\pi i\xi x}d\xi$
    $=\mathcal{F}^{-1}(\widehat{f})(x)=f(x).$
    Clearly, we have an almost everywhere results if the pointwise converge is also almost everywhere.
  3. 3.
    Let $u=\mathds{1}$. It is known that $\widehat{\mathds{1}}=\delta$. Therefore, $\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathds{1}))}=\delta\cdot(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})$. Thus, given $\psi\in\mathscr{S}(\mathbb{R})$,
    $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathds{1}))},\psi\rangle$ $=\langle\delta\cdot(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1}),\psi\rangle$
    $=\langle\delta,(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\psi\rangle=(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})(0)\psi(0).$
    Since $(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\to 1$ pointwise as $n\to\infty$,
    $\lim_{n\to\infty}\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathds{1}))},\psi\rangle=\langle\delta,\psi\rangle,$
    and so $\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathds{1}))}\to\delta=\widehat{\mathds{1}}$ as $n\to\infty$ in the weak-* topology of $\mathscr{S}^{\prime}$, i.e., in the sense of distributions. Since the Fourier transform is a homeomorphism in $\mathscr{S}^{\prime}$ it follows that $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathds{1}))\to\mathds{1}$ as $n\to\infty$ in the sense of distributions. The case for $u=\delta$ follows by noting that
    $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\delta))},\psi\rangle=\langle\mathds{1},(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\psi\rangle=\int_{\mathbb{R}}(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\psi,\quad\forall\psi\in\mathscr{S}(\mathbb{R}),$
    and then applying the LDCT. For the principal value distribution, $\widehat{\mathrm{pv}\frac{1}{x}}=-i\pi\operatorname{sign}$ as a distribution, then, given $\psi\in\mathscr{S}(\mathbb{R})$,
    $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathrm{pv}\frac{1}{x}))},\psi\rangle$ $=\langle-i\pi\operatorname{sign},(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\psi\rangle$
    $=-i\pi\int_{\mathbb{R}}\operatorname{sign}(\xi)(1-(1-\widehat{\varphi_{\varepsilon}}(\xi))^{n+1})\psi(\xi)d\xi,$
    and the result follows by applying again the LDCT. Since $\widehat{e^{2\pi i\operatorname{id}y}}=\delta_{y}$, i.e., the Dirac delta centered at $y\in\mathbb{R}$ the results follows from the previous proof. The convergence of the linear combination and translations of these distributions follow immediately.
  4. 4.
    First we prove that for the claims to hold, it is sufficient to prove that for any $\psi\in\mathscr{S}(\mathbb{R})$ then $\psi(1-\widehat{\varphi_{\varepsilon}})^{n}\to 0$ as $n\to\infty$ in the Schwartz topology. Indeed, let $f\in\mathscr{S}(\mathbb{R})$ and suppose it holds that $\widehat{f}(1-\widehat{\varphi_{\varepsilon}})^{n}\to 0$ in the Schwartz topology as $n\to\infty$. Since $\widehat{f}\in\mathscr{S}(\mathbb{R})$ and applying the same steps as in the previous proofs,
    $\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))}=\hat{f}\left(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1}\right)\to\hat{f},\quad n\to\infty$
    in the Schwartz topology by assumption. Since the Fourier transform is a homeomorphism on $\mathscr{S}(\mathbb{R})$, then
[…]
    $=\lim_{n\to\infty}\mathcal{F}^{-1}\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)=\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)),$
    in the Schwartz topology. If $f\in\mathscr{S}^{\prime}(\mathbb{R})$, since $\varphi$ is even,
    $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))},\psi\rangle=\langle T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)),\widehat{\psi}\rangle=\langle f,T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\widehat{\psi}))\rangle,\quad\forall\psi\in\mathscr{S}(\mathbb{R}).$
    Since $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\widehat{\psi}))\to\widehat{\psi}$ as $n\to\infty$ in the Schwartz topology, and tempered distributions are continuous functionals on $\mathscr{S}(\mathbb{R})$, then $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))},\psi\rangle\to\langle f,\widehat{\psi}\rangle=\langle\widehat{f},\psi\rangle$ as $n\to\infty$ in the weak-* topology, i.e., in the tempered distribution sense. Since the Fourier transform is a homeomorphism on $\mathscr{S}^{\prime}(\mathbb{R})$ the result follows as previously noted.
    So we proceed now to prove that for any $\psi\in\mathscr{S}(\mathbb{R})$ then $\psi(1-\widehat{\varphi_{\varepsilon}})^{n}\to 0$ as $n\to\infty$ in the Schwartz topology if $\varphi$ is the Gaussian Kernel [^2] Note we cannot prove that $1-(1-\widehat{\varphi_{\varepsilon}})^{n}\to 1$ in the Schwartz topology since the constant function $1$ is not in $\mathscr{S}(\mathbb{R})$, and $\mathscr{S}(\mathbb{R})$ is complete.. Assume without loss of generality that $\varepsilon=1$. Then $\widehat{\varphi}(\xi)=e^{-\xi^{2}}$ for all $\xi\in\mathbb{R}$. Let $h(x)=1-e^{-x^{2}}$ for all $x\in\mathbb{R}$. We need to prove that for all $\alpha,\beta\in\mathbb{N}_{0}$,
    $\lim_{n\to\infty}\sup_{x\in\mathbb{R}}\left|x^{\alpha}D^{\beta}(\psi h^{n})(x)\right|=0,$ (10)
    where $D$ is the differential operator, i.e., $D^{\beta}f=f^{(\beta)}$.
    Fix $\alpha,\beta\in\mathbb{N}_{0}$. For simplicity of notation, we will denote both the function $f$ and its value at $x$ as $f(x)$. First, it is not difficult to check by induction that for any $m\in\mathbb{N}$, it holds $D^{m}h(x)=P(m,x)e^{-x^{2}}$ where $P(m,x)$ is a polynomial of degree up to $m$. Applying Faa di Bruno’s formula [roman1980formula] for $m\leq n$, it holds that
    $D^{m}(h(x))^{n}=\sum\frac{m!}{k_{1}!1!^{k_{1}}k_{2}!2^{k_{2}}\dots k_{m}!m!^{k_{m}}}\frac{n!}{(n-\bar{k})!}(h(x))^{n-\bar{k}}\prod_{j=1}^{m}(D^{j}h(x))^{k_{j}},$ (11)
    where the sum is taken over all tuples $(k_{1},\dots,k_{m})$ such that $\sum_{j=1}^{m}jk_{j}=m$ and $\bar{k}:=\sum_{j=1}^{m}k_{j}$. We define for any $m\in\mathbb{N}$, $\pi(m):=\frac{m!}{k_{1}!1!^{k_{1}}k_{2}!2^{k_{2}}\dots k_{m}!m!^{k_{m}}}$. Note that $D^{j}h(x)=P(j,x)e^{-x^{2}}$ and therefore, by the properties of the exponential, $\prod_{j=1}^{m}(D^{j}h(x))^{k_{j}}=P(m!,x)e^{-\bar{k}x^{2}}$. Thus, we can write Equation (11) as
    $D^{m}(h(x))^{n}=\sum\pi(m)\frac{n!}{(n-\bar{k})!}\left(1-e^{-x^{2}}\right)^{n-\bar{k}}P(m!,k)e^{-\bar{k}x^{2}},$ (12)
    and note that the only dependence on $n$ occurs in $\frac{n!}{(n-\bar{k})!}(1-e^{-x^{2}})^{n-\bar{k}}$. Also note that $\frac{n!}{(n-\bar{k})!}\leq n^{\bar{k}}$ for all $n,\bar{k}\in\mathbb{N}$ such that $n\geq\bar{k}$, which we are going to assume from now on. Applying the Leibniz formula of differentiation to Equation (10) and inserting Equation (12), we get
    $\sup_{x\in\mathbb{R}}\left|x^{\alpha}D^{\beta}(\psi h^{n})(x)\right|$
    $\leq\sum_{m=0}^{\beta}\sum\binom{\beta}{m}\pi(m)\sup_{x\in\mathbb{R}}\left|x^{\alpha}D^{\beta-m}\psi(x)\frac{n!}{(n-\bar{k})!}(1-e^{-x^{2}})^{n-\bar{k}}P(m!,x)e^{-\bar{k}x^{2}}\right|,$
    since the sums are finite and do not depend on $n$, we just need to show that for any $m\in\{0,\dots,\beta\}$ it holds
    $\lim_{n\to\infty}\sup_{x\in\mathbb{R}}E_{n,m}(x)=0,$
    where
    $E_{n,m}(x):=\left|x^{\alpha}D^{\beta-m}\psi(x)\frac{n!}{(n-\bar{k})!}(1-e^{-x^{2}})^{n-\bar{k}}P(m!,x)e^{-\bar{k}x^{2}}\right|.$
    It is clear that $E_{n,m}(0)=0$. We proceed by breaking $\mathbb{R}$ in three intervals. The first $|x|\leq\delta$ for any $\delta<1$, the second $\delta<|x|\leq R$ for some $R\in\mathbb{R}$ and the third $|x|>R$. We will show that neither $\delta$ nor $R$ depend on $n$. We know that, by Taylor’s formula, for a sufficiently small $\delta$, it holds that $1-e^{-x^{2}}\leq x^{2}$ for all $|x|\leq\delta$. Being $x^{2}$ strictly increasing, $x^{2}\leq\delta^{2}$. Moreover, $\psi\in\mathscr{S}(\mathbb{R})$ which implies
    $C_{\alpha,\beta,m}:=\sup_{x\in\mathbb{R}}\left|x^{\alpha}D^{\beta-m}\psi(x)P(m!,x)e^{-\bar{k}x^{2}}\right|<\infty.$
    Thus
    $\sup_{|x|\leq\delta}E_{n,m}(x)\leq C_{\alpha,\beta,m}\frac{n!}{(n-\bar{k})!}\delta^{2(n-\bar{k})}\leq C_{\alpha,\beta,m}n^{\bar{k}}\delta^{2(n-\bar{k})}.$
    But the exponential decrease dominates polynomial increase, i.e., $\lim_{n\to\infty}n^{\bar{k}}\delta^{2n}=0$, and so
    $\lim_{n\to\infty}\sup_{|x|\leq\delta}E_{n,m}(x)=0.$
    Fix now $R\in\mathbb{R}$. Then, for $\delta\leq|x|\leq R$ it holds that $0<h(x)=1-e^{-x^{2}}<1$. Take $\gamma=\max_{\delta\leq|x|\leq R}h(x)<1$, which clearly exists because $h$ is even and monotonically increasing for $x\geq 0$. Then we arrive at the same conclusions as above by the same argument. Thus $\lim_{n\to\infty}\sup_{\delta<|x|\leq R}E_{n,m}(x)=0$. So we just need to prove that $R$ does not depend on $n$, because if it does not, being $\psi\in\mathscr{S}(\mathbb{R})$ and $1-h^{n}\in\mathscr{S}(\mathbb{R})$ then $\psi(1-h^{n})\in\mathscr{S}(\mathbb{R})$ and so $\psi h^{n}\in\mathscr{S}(\mathbb{R})$, which implies that $\lim_{|x|\to\infty}E_{n,m}(x)=0$, but the limit depends on $n$, so if we make it independent of $n$ we are done by splitting the domain as previously noted. For this purpose, note that being $\psi\in\mathscr{S}(\mathbb{R})$ it holds that for any $m\in\mathbb{N}_{0}$,
    $\lim_{|x|\to\infty}\left|x^{\alpha}D^{\beta-m}\psi(x)P(m!,x)\right|=0.$
    Consider the function $g_{x}(n):=\frac{n!}{(n-\bar{k})!}(1-e^{-x^{2}})^{n-\bar{k}}e^{-\bar{k}x^{2}}\geq 0$ for all $n\in\mathbb{N}$ with $n\geq\bar{k}$ and all $x\in\mathbb{R}$. Observe that all terms of $g$ are strictly positive, that $\frac{n!}{(n-\bar{k})!}\leq n^{\bar{k}}$, and also $(1-e^{x^{2}})^{n-\bar{k}}\leq e^{-(n-\bar{k})e^{-x^{2}}}$ because $1-x\leq e^{-x}$ for all $x\in[0,1]$, and $e^{-x^{2}}\in[0,1]$. Therefore, $g_{x}(n)\leq v_{x}(n):=n^{\bar{k}}e^{-(n-\bar{k})e^{-x^{2}}}e^{-\bar{k}x^{2}}$. But by calculus it is not difficult to check that $v_{x}$ has a global maximum at $n=\bar{k}e^{x^{2}}$, which implies that
    $g_{x}(n)\leq\bar{k}^{\bar{k}}e^{-\bar{k}}e^{\bar{k}e^{-x^{2}}},\quad\forall n\geq\bar{k}.$
    Therefore, the limit as $|x|\to\infty$ of $g_{x}(n)$ exists and is less than $\bar{k}^{\bar{k}}e^{-\bar{k}}$ for all $n\in\mathbb{N}$ with $n\geq\bar{k}$. Thus,
    $0\leq\lim_{|x|\to\infty}E_{n,m}(x)\leq\lim_{|x|\to\infty}\left|x^{\alpha}D^{\beta-m}\psi(x)P(m!,x)\right|\bar{k}^{\bar{k}}e^{-\bar{k}}e^{-\bar{k}e^{-x^{2}}}=0\cdot\bar{k}^{\bar{k}}e^{-\bar{k}}=0.$
    That is, given $\eta>0$ we can choose $R=R(\eta)>0$, $\delta>0$ and $N=N(\eta,\alpha,\beta,m)\in\mathbb{N}$ such that for all $n\geq N$,
    $\sup_{x\in\mathbb{R}}E_{n,m}(x)\leq\sup_{|x|\leq\delta}E_{n,m}(x)+\sup_{\delta<|x|\leq R}E_{n,m}(x)+\sup_{|x|>R}E_{n,m}(x)<\eta.$
    This finishes the proof of the Theorem.
We want to remark that the last statement of Theorem 3 is a new formula for the inverse of the Weierstrass Transform, i.e., the operator $\lim_{n\to\infty}T_{\varepsilon,n}^{-1}$ is the inverse for the Weierstrass Transform with parameter $\varepsilon>0$ for any $f\in\mathscr{S}(\mathbb{R})$ or $f\in\mathscr{S}^{\prime}(\mathbb{R})$. Note that $L^{p}(\mathbb{R})$ is injected in $\mathscr{S}^{\prime}(\mathbb{R})$ for any $p\in[1,\infty]$, as well as any $v\in L^{1}_{loc}(\mathbb{R})$ with $|v(x)|\leq C(1+|x|^{2})^{N/2}$ for some $N\in\mathbb{N}$ and $C>0$. Also note that the Delta, Heaviside and Principal Value distributions are in $\mathscr{S}^{\prime}(\mathbb{R})$ and so any distribution with compact support. Therefore, Theorem 3 presents a result that can be used in several situations which is numerically simple to compute. To end this section, note that if $f\in\mathscr{S}(\mathbb{R})$ of $f\in\mathscr{S}^{\prime}(\mathbb{R})$ and $\varphi$ is the Gaussian Kernel, then, in their respective topologies
$f=\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$ $=\lim_{n\to\infty}\sum_{k=0}^{n}(-1)^{k}\binom{n+1}{k+1}T_{\varepsilon}^{k+1}(f)$
$=\lim_{n\to\infty}\sum_{k=1}^{n+1}(-1)^{k+1}\binom{n+1}{k}T_{\varepsilon}^{k}(f)$
$=\lim_{n\to\infty}\sum_{k=0}^{n+1}(-1)^{k+1}\binom{n+1}{k}T_{\varepsilon}^{k}(f)-f,$
and so $\lim_{n\to\infty}\sum_{k=0}^{n+1}(-1)^{k+1}\binom{n+1}{k}T_{\varepsilon}^{k}(f)=\lim_{n\to\infty}-(\operatorname{id}-T_{\varepsilon})^{n+1}(f)=0$ in their respective topologies for any $f\in\mathscr{S}(\mathbb{R})$ or $f\in\mathscr{S}^{\prime}(\mathbb{R})$, and so $\lim_{n\to\infty}(\operatorname{id}-T_{\varepsilon})^{n}\equiv 0$, i.e., the zero operator on $\mathscr{S}(\mathbb{R})$ or $\mathscr{S}^{\prime}(\mathbb{R})$. Note the notation $(\operatorname{id}-T_{\varepsilon})^{n}$ is the $n$-th folding composition of $\operatorname{id}-T_{\varepsilon}$.
In Section 3 we show the numerical validation of Theorem 3.
2.3 Discussion about polynomials in the Euclidean space
This small section discusses how most of the results polynomials in the real line extend to polynomials in the Euclidean space. Note that the Binomial Theorem, as well as the Pascal Triangle and the proofs by induction extend nicely if we use multi-index notation. Indeed, let $d\in\mathbb{N}$ and $\alpha,\beta\in\mathbb{N}_{0}^{d}$, and $x,y\in\mathbb{R}^{d}$. If we define
$x^{\alpha}:=x_{1}^{\alpha_{1}}\dots x_{d}^{\alpha_{d}},\quad$ $\alpha\pm\beta:=(\alpha_{1}\pm\beta_{1},\dots,\alpha_{d}\pm\beta_{d})$
$\alpha\leq\beta\iff\alpha_{i}\leq\beta_{i},\forall i\in\{1,\dots,d\},\quad$ $\alpha!:=\alpha_{1}!\cdot\alpha_{2}!\dots\alpha_{d}!$
$\binom{\alpha}{\beta}:=\frac{\alpha!}{\beta!(\alpha-\beta)!}\quad$ $|\alpha|:=|\alpha_{1}+\dots+\alpha_{d}|,$
then it holds that
$(x\pm y)^{\alpha}=\sum_{\nu\leq\alpha}(\pm 1)^{|\alpha-\nu|}\binom{\alpha}{\nu}x^{\nu}y^{\alpha-\nu}.$
We define $\mathscr{P}_{n}(\mathbb{R}^{d})$ as the vector space of $n\in\mathbb{N}_{0}$ order polynomials in $\mathbb{R}^{d}$, that is,
$p\in\mathscr{P}_{n}(\mathbb{R}^{d})\iff p(x)=\sum_{|\alpha|\leq n}a_{\alpha}x^{\alpha},\quad\forall x\in\mathbb{R}^{d}.$
Note that a similar result to Definition 1 holds in $\mathbb{R}^{d}$. Indeed, let $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ be even. Then
$\int_{\mathbb{R}^{d}}x^{\alpha}\varphi_{\varepsilon}(x)dx=\varepsilon^{d|\alpha|}\int_{\mathbb{R}^{d}}x^{\alpha}\varphi(x)dx=\begin{cases}\varepsilon^{d|\alpha|}c_{\alpha}&|\alpha|\text{ is even or zero}\\ 0&|\alpha|\text{ is odd}\end{cases},$
where $c_{\alpha}:=\int_{\mathbb{R}^{d}}x^{n}\varphi(x)dx$, and clearly $c_{0}=1$.
Fix then any even $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ and $\varepsilon>0$. In either case, observe it is in fact immediate to generalize certain previous results. Indeed, the proof of Theorem 1 is extrapolated as is to $\mathbb{R}^{d}$, from which we get that
$(p*\varphi_{\varepsilon})(x)=\sum_{\alpha\leq n}\sum_{\begin{subarray}{c}\beta\leq\alpha\\ |\alpha-\beta|\in 2\mathbb{N}_{0}\end{subarray}}\binom{\alpha}{\beta}a_{\alpha}x^{\beta}\varepsilon^{d|\alpha-\beta|}c_{\alpha-\beta},\quad(x\in\mathbb{R}^{d}).$
That is, a similar result to Equation (3). Clearly, then
$(p*\varphi_{\varepsilon})(x)$ $=\sum_{\alpha\leq n}a_{\alpha}x^{\alpha}+\sum_{\alpha\leq n}\sum_{\begin{subarray}{c}\beta<\alpha\\ |\alpha-\beta|\in 2\mathbb{N}\end{subarray}}\binom{\alpha}{\beta}a_{\alpha}x^{\beta}\varepsilon^{d|\alpha-\beta|}c_{\alpha-\beta}$
$=:p(x)+p_{1}(x),\quad(x\in\mathbb{R}^{d}).$
which is similar to the results of equation (4) in Corollary 1. That is, the convolution of an $n$ degree polynomial in $\mathbb{R}^{d}$ with even Schwartz convolution kernel, results in the sum of the polynomial itself with a polynomial of degree n2n-2. This therefore implies that the results of Corollary 2 can be extrapolated, by noting that $|\alpha|=1\iff\alpha\in\{(1,\dots,0),(0,1\dots,0),\dots,(0,\dots,1)\}$, and therefore $x\mapsto x^{\alpha}=x_{i}$ for some $i\in\{1,\dots,d\}$. Thus, $\int_{\mathbb{R}^{d}}(x-y)^{\alpha}\varphi_{\varepsilon}(y)dy=x_{i}-\int_{\mathbb{R}^{d}}x_{i}\varphi_{\varepsilon}(y)dy=s_{i}$ for some $i\in\{1,\dots,d\}$ and all $x\in\mathbb{R}^{d}$. Thus if $p\in\mathscr{P}_{1}(\mathbb{R}^{d})$ we get
$(p*\varphi_{\varepsilon})(x)=\sum_{|\alpha|\leq 1}c_{\alpha}\int_{\mathbb{R}^{d}}(x-y)^{\alpha}\varphi_{\varepsilon}(y)dy=\sum_{|\alpha|\leq 1}c_{\alpha}x^{\alpha}=p(x),\quad(x\in\mathbb{R}^{d}),$
that is, polynomials in $\mathscr{P}_{1}(\mathbb{R}^{d})$ and $\mathscr{P}_{0}(\mathbb{R}^{d})$ are invariant to the convolution with even Schwartz convolution kernel. Thus, we can reach the same conclusions as in Theorem 2 with an identical proof. That is, convolution with an even convolution kernel is an isomorphism of $\mathscr{P}_{n}(\mathbb{R}^{d})$ onto itself for all $n\in\mathbb{N}_{0}$, and we can compute the inverse of the convolution similar to equations (7) and (8). Indeed, if $n$ is even the inverse of $p\in\mathscr{P}_{n}(\mathbb{R}^{d})\mapsto T_{\varepsilon}(p)=p*\varphi_{\varepsilon}\in C^{\infty}(\mathbb{R}^{d})$ with $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ even is
$T_{\varepsilon}^{-1}(p)=\sum_{|\alpha|\leq n/2}\sum_{\beta\leq\alpha}(-1)^{|\alpha+\beta|}\binom{\alpha}{\beta}T_{\varepsilon}^{|\alpha-\beta|}(p),\quad(p\in\mathscr{P}_{n}(\mathbb{R}^{d})),$
while if $n$ is odd
$T_{\varepsilon}^{-1}(p)=\sum_{|\alpha|\leq\frac{n-1}{2}}\sum_{\beta\leq\alpha}(-1)^{|\alpha+\beta|}\binom{\alpha}{\beta}T_{\varepsilon}^{|\alpha-\beta|}(p),\quad(p\in\mathscr{P}_{n}(\mathbb{R}^{d})).$
3 Applications
In this section we present two particular applications of the results in Section 2. First we validate numerically Theorem 2, and then we apply the results of Theorem 3 to an application in the reconstruction of a sinusoidal function.
Figure 1: Representation of a numerical validation of Theorem 2. The left picture represents as a solid black line the original polynomial $p$ which is the $n=50$ order Taylor polynomial of the function $x\in\mathbb{R}\mapsto\sin(5x)+\sin(3x)$. It also represents as a blue solid line its convolution $p*\varphi_{\varepsilon}$ with convolution kernel $s\mapsto\varphi(x)=\operatorname{ind}_{(-1,1)}\exp(-1/(1-x^{2}))$ and parameter $\varepsilon=0.9$. The dashed red line represents $T_{\varepsilon}^{-1}(p*\varphi_{\varepsilon})$ computed using equation (7). The right picture represents the error between the original polynomial and the inverse of the convolution.
3.1 Validation of Theorem 2
In order to numerically validate Theorem 2, we present Figure 1. The polynomial is an $n=50$ Taylor polynomial of the function $x\in\mathbb{R}\mapsto\sin(3x)+\sin(5x)$. Due to the fact that, numerically speaking, the domain of the original polynomial is always bounded, it can be seen that at the ends of the domain the inverse of the convolution and the original polynomial differs the most. This is a fact on how convolution is computed numerically. Due to being the functions actually finite sequences, a zero padding is needed in order to be able to compute the convolution numerically. Then, once computed, we take back the same number of samples as the original sequence in order to compare it with the original function in its domain of definition. Nevertheless, it can be seen that, once the effect of the edges is negligible, that Figure 1 validates numerically Theorem 2, recovering $p$ from $p*\varphi_{\varepsilon}$ by applying Equation (7). Note that the used convolution kernel has compact support and zeros in its Fourier transform (cf. Theorem 3).
3.2 Applications to Signal Processing
We are going to consider the problem of deconvolution of an arbitrary function. Suppose that the convolution kernel is the Gaussian Kernel, $\varphi\in\mathscr{S}(\mathbb{R})$. Suppose then, that some function $f:\mathbb{R}\to\mathbb{R}$ has been convolved with $\varphi_{\varepsilon}$ for some $\varepsilon>0$. Suppose further that $f$ satisfies any condition of Theorem 3, Then, by letting $n\to\infty$ in Equation (7) we could recover $f$ (in the appropriate convergence).
The traditional approach to deconvolution consists on working on the Fourier transformed space. For example, under the appropriate conditions, for $\varphi$ the Gaussian Kernel and $f\in\mathscr{S}(\mathbb{R})$ it holds that $\widehat{f*\varphi}(\xi)=\widehat{f}\widehat{\varphi}(\xi)\Longrightarrow f=\mathcal{F}^{-1}\left(\widehat{f*\varphi}e^{\operatorname{id}^{2}}\right)$. Note this is ill-posed numerically speaking, since $x\mapsto e^{x^{2}}$ growths exponentially fast, and in particular, this function does not posses inverse Fourier transform. However, by using the results of Theorem 3, we can reconstruct $f$ directly from the samples of $f*\varphi$, without the need of passing to the frequency domain.
Consider then, the function $t\in\mathbb{R}\mapsto f(t)=\sin(5t)+\sin(3t)$ (we denote $t$ for time) as in Figure 1 and let $\varphi$ be the Gaussian convolution kernel. Since $f$ defines a regular tempered distribution, the assumptions of Theorem 3 hold.
Figure 2: Representation of the numerical validation of Theorem 3 under ideal conditions. Left picture represents in black the original function $t\mapsto f(t)=\sin(5t)+\sin(3t)$, in blue its convolution $T_{\varepsilon}(f)$ with the Gaussian Kernel $\varphi$ with $\varepsilon=0.55$, and as dashed red line its reconstruction $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$ using Equation (9) for $n=90$. Right picture represents in black the Fast Fourier Transform (FFT) of $f$, in green the FFT of $\varphi$, in blue the FFT of $T_{\varepsilon}(f)$ and in red the FFT of the reconstructed function $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$.
Note $T_{\varepsilon}(f)=f*\varphi_{\varepsilon}$ for $\varepsilon>0$ is the available data. In Figure 2 we can see the results of applying Theorem 3 to $T_{\varepsilon}(f)$. As it can be seen, the signal reconstruction is almost perfect. Theoretically speaking, we would need to let $n\to\infty$ for (distributional) convergence, but at $n\geq 100$ the numerical errors start to affect the method, and the recovery of the signal for $n=90$ is sufficiently good, numerically speaking, as can be seen from the time and frequency domains. Indeed, by looking at the spectra of $f$, $T_{\varepsilon}(f)$ and $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$, it is well validated that the method works, since we were capable of recovering (numerically speaking) the original function.
Now we consider the situation depicted in Figure 3. We suppose that $T_{\varepsilon}(f)$ has been modified with additive noise $\eta:\mathbb{R}\to\mathbb{R}$ with zero mean, and variance $0.5$. Note that the signal to noise ratio is less than one, as it can be seen from the top and bottom left pictures. However, by applying Equation (9) to $T_{\varepsilon}(f)+\eta$, it can be seen from the top right picture that we still recover (with amplified noise) the frequency contents that were once removed by the convolution kernel. This can be “heuristically” be justified, by noting that if $T_{\varepsilon,n}^{-1}$ is the inverse (in the limit) of the convolution of “sufficiently nice” functions, then this operator must amplify noise, being a deconvolution. Thus, after computing Equation (9) it is clear that the original function was (at least) composed of higher frequencies than the spectrum of $T_{\varepsilon}(f)+\eta$ shows. Thus, we can apply a low pass filter $h:\mathbb{R}\to\mathbb{R}$ to the recovered noisy signal in order to remove the undesired noise. This is shown in the bottom pictures. As it can be seen, we have effectively recovered a close version to the original signal, both in the time and frequency domain, even if noise was present, with the particularity that the signal to noise ratio was lower than one.
Figure 3: Representation of the numerical validation of Theorem 3 under non ideal conditions. Top and bottom left pictures represent in black the original function $t\mapsto f(t)=\sin(5t)+\sin(3t)$, with a blue line $T_{\varepsilon}(f)$ with $\varepsilon=0.55$ and $\varphi$ the Gaussian Kernel, as a yellow line $T_{\varepsilon}(f)+\eta$, where $\eta$ is noise generated using a normal distribution with zero mean and variance 0, and with a red line $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)$ where $n=90$. Top right picture depicts as a black line the FFT of $f$, as a yellow line the FFT of $T_{\varepsilon}(f)+\eta$ and as a red line $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)$, i.e., the application of Equation (9) to $T_{\varepsilon}(f)+\eta$. The bottom right picture shows with a black line the FFT of $f$, with a green line the FFT of the used filter $t\mapsto h(t)=2\operatorname{sinc}(2t)$, with a yellow line the FFT of $T_{\varepsilon}(f)+\eta$ and with a red line the FFT of $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)*h$, that is, $g:=T_{\varepsilon}(f)+\eta$.
4 Conclusions
In this work we have shown that even Schwartz convolution (i.e., convolution with even Schwartz convolution kernels) is an automorphism between polynomial spaces of finite order. In particular, we obtained an explicit formula for the inversion of the convolution, and numerically validated it with the Taylor polynomial of a sinusoidal function. Moreover, we have shown how these results can be extended to a bigger class of functions and tempered distributions, having a new iterative formula for the inverse of the convolution of these functions and distributions with Schwartz kernels. We have numerically validated this extension with an application to signal processing in which a sinusoidal function has been recovered after convolution, either in a noiseless scenario and in a less than unity signal to noise ratio situation.
5 Acknowledgments
This work has been supported by the FPU program of the Ministry of science, innovation and universities of Spain. The author thanks Juan F. Jiménez for his helpful feedback that contributed to the refinement of this paper.